# presep= ]e1,e2,e3

### v1

@@ Appendix B Proofs of results from Section 4
$\log(u_{n})=\frac{1}{\alpha_{x}}\log\left(\ell^{x}(u_{n})\right)-\frac{1}{\alpha_{x}}\log\left(\frac{k_{n}}{n}\right).$
@@
$h_{n}\log(u_{n})$ $=h_{n}\left(\frac{1}{\alpha_{x}}\log\left(\ell^{x}(u_{n})\right)-\frac{1}{\alpha_{x}}\log\left(\frac{k_{n}}{n}\right)\right)$
$=(\overline{F}^{x}(u_{n}))^{e}\left(\frac{1}{\alpha_{x}}\log\left(\ell^{x}(u_{n})\right)-\frac{1}{\alpha_{x}}\log\left(n^{-(1-\delta)}\right)\right)$
$=\alpha_{x}^{-1}u_{n}^{-\alpha_{x}e}\ell^{x}(u_{n})^{e}\log\left(\ell^{x}(u_{n})\right)+\mathcal{O}\left(n^{-(1-\delta)e}\log(n)\right)$
$\sim\alpha_{x}^{-1}u_{n}^{-\alpha_{x}e}\ell^{x}(u_{n})^{e}\log\left(\ell^{x}(u_{n})\right)$
$\leq\alpha_{x}^{-1}u_{n}^{-\alpha_{x}e}\ell^{x}(u_{n})^{1+e}=\alpha_{x}^{-1}\left(u_{n}^{-\alpha_{x}\left(\frac{e}{1+e}\right)}\ell^{x}(u_{n})\right)^{1+e}\to 0.$
Next
$nh_{n}\overline{F}^{x}(u_{n})\asymp n\left(\overline{F}^{x}(u_{n})\right)^{1+e}\asymp n\left(\frac{k_{n}}{n}\right)^{1+e}=n^{-(1-\delta)(1+e)+1}=n^{-(1-\delta)e+\delta}\to\infty.$
As for the third condition we have
$\frac{\overline{F}^{x}(u_{n})+h_{n}^{c_{x}\eta-2}}{nh_{n}^{c_{x}}(\overline{F}^{x}(u_{n}))^{2}}=\frac{1}{nh_{n}^{c_{x}}\overline{F}^{x}(u_{n})}+\frac{h_{n}^{c_{x}(\eta-1)-2}}{n(\overline{F}^{x}(u_{n}))^{2}}.$
The first term vanishes asymptotically since
$\left(nh_{n}^{c_{x}}\overline{F}^{x}(u_{n})\right)^{-1}$ $\asymp\left(n(\overline{F}^{x}(u_{n}))^{c_{x}e+1}\right)^{-1}=\left(n\left(\frac{k_{n}}{n}\right)^{c_{x}e+1}\right)^{-1}\asymp\left(n^{1-(1-\delta)(c_{x}e+1)}\right)^{-1}\to 0.$
Moreover
$\frac{h_{n}^{c_{x}(\eta-1)-2}}{n(\overline{F}^{x}(u_{n}))^{2}}=\frac{1}{nh_{n}\overline{F}^{x}(u_{n})}\frac{h_{n}^{c_{x}(\eta-1)-1}}{\overline{F}^{x}(u_{n})}\asymp n^{(1-\delta)e-\delta}\overline{F}^{x}(u_{n})^{e(c_{x}(\eta-1)-1)-1},$
which vanishes as $n\to\infty$, since $e>(c_{x}(\eta-1)-1)^{-1}$. ∎
Proof of Theorem 7.
Note that $nh_{n}\overline{F}^{x}(u_{n})=k_{n}h_{n}\asymp\log(n)^{2-e}\to\infty$. Now, the relation
$\log\left(\frac{n}{k_{n}}\right)=-\log\left(\overline{F}^{x}(u_{n})\right)=\alpha_{x}\log(u_{n})-\log(\ell^{x}(u_{n})),$
implies that
$\frac{\log(n/k_{n})}{\alpha_{x}\log(u_{n})}=1-\frac{\log(\ell^{x}(u_{n}))}{\alpha_{x}\log(u_{n})}\to 1,$
using Bingham, Goldie and Teugels (1987) Proposition 1.3.6 (i). Hence $\log\left(n/k_{n}\right)\sim\alpha_{x}\log(u_{n})$ and consequently
$h_{n}\log(u_{n})\sim\frac{h_{n}\log(n/k_{n})}{\alpha_{x}}\leq\frac{h_{n}\log(n)}{\alpha_{x}}\asymp\frac{\log(n)^{1-e}}{\alpha_{x}}\to 0.$
@@
By definition of $u_{n}$ and properties of generalised inverses, it holds that $n\overline{F}^{x}(u_{n})=k_{n}$. Assumption (7) implies that $u_{n}=\mathcal{O}((n/k_{n})^{1/\alpha_{x}})$.
Then recall that
$B_{n}^{x}(s_{0})=\sup_{t\geq s_{0}}\left|\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-\alpha_{x}}g(x)\right|.$
The above difference can be split as follows
$\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-\alpha_{x}}g(x)=\int K\left(u\right)\left(\frac{\overline{F}^{x-uh_{n}}(tu_{n})}{\overline{F}^{x}(u_{n})}g(x-h_{n}u)-t^{-\alpha_{x}}g(x)\right)\mathrm{d}u$
$=\int K\left(u\right)t^{-\alpha_{x}}\left(g(x-h_{n}u)-g(x)\right)\mathrm{d}u$
$\quad+\int K\left(u\right)\left(\frac{\overline{F}^{x-uh_{n}}(tu_{n})}{\overline{F}^{x}(u_{n})}-\frac{\overline{F}^{x}(tu_{n})}{\overline{F}^{x}(u_{n})}\right)g(x-h_{n}u)\mathrm{d}u+\int K\left(u\right)\left(\frac{\overline{F}^{x}(tu_{n})}{\overline{F}^{x}(u_{n})}-t^{-\alpha_{x}}\right)g(x-h_{n}u)\mathrm{d}u.$
$\int K\left(u\right)t^{-\alpha_{x}}\left(g(x-h_{n}u)-g(x)\right)\mathrm{d}u\leq t^{-\alpha_{x}}h_{n}c_{g}\int K\left(u\right)\left|u\right|\mathrm{d}u=\mathcal{O}(t^{-\alpha_{x}}h_{n}).$
When examining the second term, we rely on calculations from the proof of Lemma 4 and utilize the inequality $\left|e^{z}-1\right|\leq\left|z\right|+(1+e^{\left|z\right|})w^{2}/2$. This yields
$\left(\frac{\overline{F}^{x-uh_{n}}(tu_{n})}{\overline{F}^{x}(u_{n})}-\frac{\overline{F}^{x}(tu_{n})}{\overline{F}^{x}(u_{n})}\right)\leq\frac{\overline{F}^{x}(tu_{n})}{\overline{F}^{x}(u_{n})}\left|\frac{\overline{F}^{x-uh_{n}}(tu_{n})}{\overline{F}^{x}(tu_{n})}-1\right|$
$\leq\frac{\overline{F}^{x}(tu_{n})}{\overline{F}^{x}(u_{n})}\left(\left|\log\left(\frac{\overline{F}^{x-uh_{n}}(tu_{n})}{\overline{F}^{x}(tu_{n})}\right)\right|+\frac{1}{2}\log\left(\frac{\overline{F}^{x-uh_{n}}(tu_{n})}{\overline{F}^{x}(tu_{n})}\right)^{2}\left(1+\frac{\overline{F}^{x-uh_{n}}(tu_{n})}{\overline{F}^{x}(tu_{n})}\right)\right)$
$\leq\text{cst}\ \frac{\overline{F}^{x}(tu_{n})}{\overline{F}^{x}(u_{n})}\log(tu_{n})\left|u\right|h_{n}\left(1+\frac{1}{2}\log(tu_{n})\left|u\right|h_{n}\left(1+(tu_{n})^{(c_{\alpha}+c_{L})\left|u\right|h_{n}}\right)\right)$
$\leq\text{cst}\ t^{-\alpha_{x}+\varepsilon}\log(tu_{n})\left|u\right|h_{n}\left(1+\frac{1}{2}\log(tu_{n})\left|u\right|h_{n}\left(1+(tu_{n})^{(c_{\alpha}+c_{L})\left|u\right|h_{n}}\right)\right),$
for $\varepsilon\in(0,\alpha_{x})$. The first term is the leading term and of order
$\mathcal{O}\left(t^{-\alpha_{x}+\varepsilon}\log(tu_{n})h_{n}\right)=\mathcal{O}\left(t^{-\alpha_{x}+\varepsilon}\log\left(t\frac{n}{k_{n}}\right)h_{n}\right).$
To determine the order at which the third and final term converge to zero the second order condition is needed. Indeed
$\left|\frac{\overline{F}^{x}(tu_{n})}{\overline{F}^{x}(u_{n})}-t^{-\alpha_{x}}\right|$
$=\left|\frac{n\overline{F}^{x}(tu_{n})}{k_{n}}-t^{-\alpha_{x}}\right|\leq\frac{n}{k_{n}}\left(\left|\overline{F}^{x}(tu_{n})-c_{x}(u_{n}t)^{-\alpha_{x}}\right|+t^{-\alpha_{x}}\left|c_{x}(u_{n})^{-\alpha_{x}}-\overline{F}^{x}(u_{n})\right|\right)$
$\leq\text{cst}\frac{n}{u_{n}^{\beta_{x}}k_{n}}\left(t^{-\beta_{x}}+t^{-\alpha_{x}}\right)\leq\text{cst}\frac{n}{u_{n}^{\beta_{x}}k_{n}}t^{-\alpha_{x}}=\mathcal{O}\left(k_{n}^{\frac{\beta_{x}}{\alpha_{x}}-1}n^{1-\frac{\beta_{x}}{\alpha_{x}}}t^{-\alpha_{x}}\right).$
$\sqrt{k_{n}h_{n}}B_{n}^{x}(s_{0})$ $\leq\mathcal{O}(\sqrt{k_{n}h_{n}}h_{n})+\mathcal{O}\left(\sqrt{k_{n}h_{n}}\log\left(\frac{n}{k_{n}}\right)h_{n}\right)+\mathcal{O}\left(\sqrt{k_{n}h_{n}}k_{n}^{\frac{\beta_{x}}{\alpha_{x}}-1}n^{1-\frac{\beta_{x}}{\alpha_{x}}}\right)$
$=\mathcal{O}\left(\sqrt{k_{n}h_{n}^{3}}\log\left(\frac{n}{k_{n}}\right)\right)+\mathcal{O}\left(\sqrt{h_{n}}k_{n}^{\frac{\beta_{x}}{\alpha_{x}}-\frac{1}{2}}n^{1-\frac{\beta_{x}}{\alpha_{x}}}\right)\to 0.$
This concludes condition 5. As for condition 7 we have
$\int_{1}^{\infty}\frac{\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-\alpha_{x}}g(x)}{t}\mathrm{d}t\leq\mathcal{O}(h_{n})\int_{1}^{\infty}t^{-\alpha_{x}-1}\mathrm{d}t+\mathcal{O}\left(k_{n}^{\frac{\beta_{x}}{\alpha_{x}}-1}n^{1-\frac{\beta_{x}}{\alpha_{x}}}\right)\int_{1}^{\infty}t^{-\alpha_{x}-1}\mathrm{d}t$
$+\mathcal{O}(h_{n})\int_{1}^{\infty}t^{-\alpha_{x}+\varepsilon-1}\log(tu_{n})\left(1+\frac{1}{2}\log(tu_{n})h_{n}\left(1+(tu_{n})^{(c_{\alpha}+c_{L})\left|u\right|h_{n}}\right)\right)\mathrm{d}t,$
for $\varepsilon\in(0,\alpha_{x})$. The first two integrals are obviously finite, and thus we examine the last integral, which itself consists of three sub-integrals. The first of these subintegrals is finite, and of order $\mathcal{O}(h_{n}\log(u_{n}))$, likewise for the second which is of order $\mathcal{O}(h_{n}^{2}\log(u_{n}))$. Finally, let $\varepsilon^{\prime}\in(0,\alpha_{x}-\varepsilon)$ be given. It is possible to find $n^{\prime}$ such that $\left(c_{\alpha}+c_{L}\right)h_{n}\leq\varepsilon^{\prime}$ for all $n\geq n^{\prime}$. For such $n$’s the third integral is also finite and of order $\mathcal{O}(h_{n}\log(u_{n}))$, since $u_{n}^{h_{n}}=\mathcal{O}(1)$.
Consequently
$\sqrt{k_{n}h_{n}}\int_{1}^{\infty}\frac{\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-\alpha_{x}}g(x)}{t}\mathrm{d}t\leq\mathcal{O}\left(\sqrt{k_{n}h_{n}^{3}}\log\left(\frac{n}{k_{n}}\right)\right)+\mathcal{O}\left(\sqrt{h_{n}}k_{n}^{\frac{\beta_{x}}{\alpha_{x}}-\frac{1}{2}}n^{1-\frac{\beta_{x}}{\alpha_{x}}}\right)\to 0.$
Proof of Lemma 2.
As above, it holds that $\overline{F}^{x}(tu_{n})/\overline{F}^{x}(u_{n})=\overline{F}^{x}(tu_{n})n/k_{n}$. Consequently
$\left|\frac{\overline{F}^{x}(tu_{n})}{\overline{F}^{x}(u_{n})}-t^{-\alpha_{x}}\right|=\frac{n}{k_{n}}\left|\overline{F}^{x}(tu_{n})-c_{x}\log(u_{n}t)(u_{n}t)^{-\alpha_{x}}+c_{x}\log(u_{n}t)(u_{n}t)^{-\alpha_{x}}-\overline{F}^{x}(u_{n})t^{-\alpha_{x}}\right|$
$\leq\frac{n}{k_{n}}\Bigg{(}\left|\overline{F}^{x}(tu_{n})-c_{x}\log(u_{n}t)(u_{n}t)^{-\alpha_{x}}\right|t^{-\alpha_{x}}\left|c_{x}\log(u_{n})u_{n}^{-\alpha_{x}}-\overline{F}^{x}(u_{n})\right|+c_{x}\log(t)(u_{n}t)^{-\alpha_{x}}\Bigg{)}$
$\leq\text{cst}\frac{n}{k_{n}}\left((tu_{n})^{-\beta_{x}}+t^{-\alpha_{x}}u_{n}^{-\beta_{x}}+c_{x}\log(t)(u_{n}t)^{-\alpha_{x}}\right)\leq\text{cst}\frac{n}{k_{n}}t^{-\alpha_{x}}u_{n}^{-\alpha_{x}}\left(1+\log(t)\right).$
@@
$u_{n}^{-\alpha_{x}}\log(u_{n})=\frac{k_{n}}{n}+\mathcal{O}\left(u_{n}^{-\beta_{x}}\right),$

### v2

@@ Appendix B Proofs of results from Section 4
Due to the fact that $\delta\in((\eta+1)^{-1},1)$, we can let
$e\in\left(0,\frac{\delta}{1-\delta}\wedge\frac{\delta(\eta+1)-1}{2(1-\delta)}\right),$
since the interval is non-empty.
@@
$\log(u_{n}^{x})=\gamma(x)\log\left(\ell^{x}(u_{n}^{x})\right)-\gamma(x)\log\left(k_{n}/n\right).$
Consequently, for any strictly positive $e$ it holds that
$h_{n}\log(u_{n}^{x})$ $=h_{n}\left(\gamma(x)\log\left(\ell^{x}(u_{n}^{x})\right)-\gamma(x)\log\left(k_{n}/n\right)\right)$
$=(\overline{F}^{x}(u_{n}^{x}))^{e}\left(\gamma(x)\log\left(\ell^{x}(u_{n}^{x})\right)-\gamma(x)\log\left(n^{-(1-\delta)}\right)\right)$
$=\gamma(x)(u_{n}^{x})^{-e/\gamma(x)}\ell^{x}(u_{n}^{x})^{e}\log\left(\ell^{x}(u_{n}^{x})\right)+\mathcal{O}\left(n^{-(1-\delta)e}\log(n)\right)$
$\sim\gamma(x)(u_{n}^{x})^{-e/\gamma(x)}\ell^{x}(u_{n}^{x})^{e}\log\left(\ell^{x}(u_{n}^{x})\right)$
$\leq\gamma(x)(u_{n}^{x})^{-e/\gamma(x)}\ell^{x}(u_{n}^{x})^{1+e}=\gamma(x)\left((u_{n}^{x})^{-(e/1+e)/\gamma(x)}\ell^{x}(u_{n}^{x})\right)^{1+e}\to 0.$
The upper bound on $e$ ensures
$nh_{n}\overline{F}^{x}(u_{n}^{x})\asymp n\left(\overline{F}^{x}(u_{n}^{x})\right)^{1+e}\asymp n\left(\frac{k_{n}}{n}\right)^{1+e}=n^{-(1-\delta)(1+e)+1}=n^{-(1-\delta)e+\delta}\to\infty.$
As for the third condition we have
$\frac{\overline{F}^{x}(u_{n}^{x})+h_{n}^{c_{x}\eta-2}}{nh_{n}^{c_{x}}(\overline{F}^{x}(u_{n}^{x}))^{2}}=\frac{1}{nh_{n}^{c_{x}}\overline{F}^{x}(u_{n}^{x})}+\frac{h_{n}^{c_{x}(\eta-1)-2}}{n(\overline{F}^{x}(u_{n}^{x}))^{2}}.$
Let $c_{x}$ be a positive constant such that
$c_{x}\in\left(\frac{2e(1-\delta)+1-2\delta}{e(1-\delta)(\eta-1)}\vee 0,\frac{\delta}{e(1-\delta)}\right),$
and note that the restrictions on $\delta$ and $e$ ensures that this interval is non-empty. The upper bound on $c_{x}$ ensures that the first of the above two terms vanishes asymptotically since
$\left(nh_{n}^{c_{x}}\overline{F}^{x}(u_{n}^{x})\right)^{-1}$ $\asymp\left(n(\overline{F}^{x}(u_{n}^{x}))^{c_{x}e+1}\right)^{-1}=\left(n\left(\frac{k_{n}}{n}\right)^{c_{x}e+1}\right)^{-1}\asymp\left(n^{1-(1-\delta)(c_{x}e+1)}\right)^{-1}\to 0,$
while the lower bound ensures
$\frac{h_{n}^{c_{x}(\eta-1)-2}}{n(\overline{F}^{x}(u_{n}^{x}))^{2}}$ $=\frac{1}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\frac{h_{n}^{c_{x}(\eta-1)-1}}{\overline{F}^{x}(u_{n}^{x})}$
$\asymp n^{(1-\delta)e-\delta}\overline{F}^{x}(u_{n}^{x})^{e(c_{x}(\eta-1)-1)-1}\asymp n^{(1-\delta)e-\delta-(1-\delta)(e(c_{x}(\eta-1)-1)-1)}\to 0,$
as $n\to\infty$. ∎
Proof of Theorem 7.
Note that $nh_{n}\overline{F}^{x}(u_{n}^{x})=k_{n}h_{n}\asymp\log(n)^{2-e}\to\infty$. Now, the relation
$\log\left(n/k_{n}\right)=-\log\left(\overline{F}^{x}(u_{n}^{x})\right)=(1/\gamma(x))\log(u_{n}^{x})-\log(\ell^{x}(u_{n}^{x})),$
implies that
$\frac{\gamma(x)\log(n/k_{n})}{\log(u_{n}^{x})}=1-\frac{\gamma(x)\log(\ell^{x}(u_{n}^{x}))}{\log(u_{n}^{x})}\to 1,$
using Bingham1987 Proposition 1.3.6 (i). Hence $\log\left(n/k_{n}\right)\sim(1/\gamma(x))\log(u_{n}^{x})$ and consequently
$h_{n}\log(u_{n}^{x})\sim\gamma(x)h_{n}\log(n/k_{n})\leq\gamma(x)h_{n}\log(n)\asymp\gamma(x)\log(n)^{1-e}\to 0.$
Proof of Lemma 1.
By definition of $u_{n}^{x}$ and properties of generalised inverses, it holds that $n\overline{F}^{x}(u_{n}^{x})=k_{n}$. The second order assumption of equation (4) implies that $u_{n}^{x}=\mathcal{O}((n/k_{n})^{\gamma(x)})$.
Then recall that
$B_{n}^{x}(s_{0})=\sup_{t\geq s_{0}}\left|\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-1/\gamma(x)}g(x)\right|.$
The above difference can be split as follows
$\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-1/\gamma(x)}g(x)=\int K\left(u\right)\left(\frac{\overline{F}^{x-uh_{n}}(tu_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}g(x-h_{n}u)-t^{-1/\gamma(x)}g(x)\right)\mathrm{d}u$
$=\int K\left(u\right)t^{-1/\gamma(x)}\left(g(x-h_{n}u)-g(x)\right)\mathrm{d}u+\int K\left(u\right)\left(\frac{\overline{F}^{x}(tu_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}-t^{-1/\gamma(x)}\right)g(x-h_{n}u)\mathrm{d}u$
$\quad+\int K\left(u\right)\left(\frac{\overline{F}^{x-uh_{n}}(tu_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}-\frac{\overline{F}^{x}(tu_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}\right)g(x-h_{n}u)\mathrm{d}u.$
$\int K\left(u\right)t^{-1/\gamma(x)}\left(g(x-h_{n}u)-g(x)\right)\mathrm{d}u\leq t^{-1/\gamma(x)}h_{n}c_{g}\int K\left(u\right)\left|u\right|\mathrm{d}u=\mathcal{O}(t^{-1/\gamma(x)}h_{n}).$
When examining the third term, we rely on calculations from the proof of Lemma 4 and utilize the inequality $\left|e^{z}-1\right|\leq\left|z\right|+(1+e^{\left|z\right|})z^{2}/2$. This yields
$\left(\frac{\overline{F}^{x-uh_{n}}(tu_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}-\frac{\overline{F}^{x}(tu_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}\right)\leq\frac{\overline{F}^{x}(tu_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}\left|\frac{\overline{F}^{x-uh_{n}}(tu_{n}^{x})}{\overline{F}^{x}(tu_{n}^{x})}-1\right|$
$\leq\frac{\overline{F}^{x}(tu_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}\left(\left|\log\left(\frac{\overline{F}^{x-uh_{n}}(tu_{n}^{x})}{\overline{F}^{x}(tu_{n}^{x})}\right)\right|+\frac{1}{2}\log\left(\frac{\overline{F}^{x-uh_{n}}(tu_{n}^{x})}{\overline{F}^{x}(tu_{n}^{x})}\right)^{2}\left(1+\frac{\overline{F}^{x-uh_{n}}(tu_{n}^{x})}{\overline{F}^{x}(tu_{n}^{x})}\right)\right)$
$\leq\text{cst}\ \frac{\overline{F}^{x}(tu_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}\log(tu_{n}^{x})\left|u\right|h_{n}\left(1+\frac{1}{2}\log(tu_{n}^{x})\left|u\right|h_{n}\left(1+(tu_{n}^{x})^{(c_{\gamma}+c_{L})\left|u\right|h_{n}}\right)\right)$
$\leq\text{cst}\ t^{-1/\gamma(x)+\varepsilon}\log(tu_{n}^{x})\left|u\right|h_{n}\left(1+\frac{1}{2}\log(tu_{n}^{x})\left|u\right|h_{n}\left(1+(tu_{n}^{x})^{(c_{\gamma}+c_{L})\left|u\right|h_{n}}\right)\right),$
for $\varepsilon\in(0,1/\gamma(x))$. The first term is the leading term and of order
$\mathcal{O}\left(t^{-1/\gamma(x)+\varepsilon}\log(tu_{n}^{x})h_{n}\right)=\mathcal{O}\left(t^{-1/\gamma(x)+\varepsilon}\log\left(t\frac{n}{k_{n}}\right)h_{n}\right).$
To determine the order at which the second and final term converge to zero the second order condition is needed. Indeed
$\left|\frac{\overline{F}^{x}(tu_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}-t^{-1/\gamma(x)}\right|=\left|\frac{n\overline{F}^{x}(tu_{n}^{x})}{k_{n}}-t^{-1/\gamma(x)}\right|$
$\leq\frac{n}{k_{n}}\left(\left|\overline{F}^{x}(tu_{n}^{x})-c_{x}(u_{n}^{x}t)^{-1/\gamma(x)}\right|+t^{-1/\gamma(x)}\left|c_{x}(u_{n}^{x})^{-1/\gamma(x)}-\overline{F}^{x}(u_{n}^{x})\right|\right)$
$\leq\text{cst}\frac{n}{(u_{n}^{x})^{\zeta_{x}}k_{n}}\left(t^{-\zeta_{x}}+t^{-1/\gamma(x)}\right)\leq\text{cst}\frac{n}{(u_{n}^{x})^{\zeta_{x}}k_{n}}t^{-1/\gamma(x)}=\mathcal{O}\left(k_{n}^{\zeta_{x}\gamma(x)-1}n^{1-\zeta_{x}\gamma(x)}t^{-1/\gamma(x)}\right).$
$\sqrt{k_{n}h_{n}}B_{n}^{x}(s_{0})$ $\leq\mathcal{O}(\sqrt{k_{n}h_{n}}h_{n})+\mathcal{O}\left(\sqrt{k_{n}h_{n}}\log\left(n/k_{n}\right)h_{n}\right)+\mathcal{O}\left(\sqrt{k_{n}h_{n}}k_{n}^{\zeta_{x}\gamma(x)-1}n^{1-\zeta_{x}\gamma(x)}\right)$
$=\mathcal{O}\left(\sqrt{k_{n}h_{n}^{3}}\log\left(n/k_{n}\right)\right)+\mathcal{O}\left(\sqrt{h_{n}}k_{n}^{\zeta_{x}\gamma(x)-\frac{1}{2}}n^{1-\zeta_{x}\gamma(x)}\right)\to 0.$
This concludes condition 5. As for condition 7 we have
$\int_{1}^{\infty}\frac{\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-1/\gamma(x)}g(x)}{t}\mathrm{d}t$
$\leq\mathcal{O}(h_{n})\int_{1}^{\infty}t^{-1/\gamma(x)-1}\mathrm{d}t+\mathcal{O}\left(k_{n}^{\zeta_{x}\gamma(x)-1}n^{1-\zeta_{x}\gamma(x)}\right)\int_{1}^{\infty}t^{-1/\gamma(x)-1}\mathrm{d}t$
$\quad+\mathcal{O}(h_{n})\int_{1}^{\infty}t^{-1/\gamma(x)+\varepsilon-1}\log(tu_{n}^{x})\left(1+\frac{1}{2}\log(tu_{n}^{x})h_{n}\left(1+(tu_{n}^{x})^{(c_{\gamma}+c_{L})\left|u\right|h_{n}}\right)\right)\mathrm{d}t,$
for $\varepsilon\in(0,1/\gamma(x))$. The first two integrals are obviously finite, and thus we examine the last integral, which itself consists of three sub-integrals. The first of these subintegrals is finite, and of order $\mathcal{O}(h_{n}\log(u_{n}^{x}))$, likewise for the second which is of order $\mathcal{O}(h_{n}^{2}\log(u_{n}^{x})^{2})$. Finally, let $\varepsilon^{\prime}\in(0,1/\gamma(x)-\varepsilon)$ be given. It is possible to find $n^{\prime}$ such that $\left(c_{\gamma}+c_{L}\right)h_{n}\leq\varepsilon^{\prime}$ for all $n\geq n^{\prime}$. For such $n$’s the third integral is also finite and of order $\mathcal{O}(h_{n}\log(u_{n}^{x}))$, since $(u_{n}^{x})^{h_{n}}=\mathcal{O}(1)$.
@@
$\sqrt{k_{n}h_{n}}\int_{1}^{\infty}\frac{\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-1/\gamma(x)}g(x)}{t}\mathrm{d}t\leq\mathcal{O}\left(\sqrt{k_{n}h_{n}^{3}}\log\left(n/k_{n}\right)\right)+\mathcal{O}\left(\sqrt{h_{n}}k_{n}^{\zeta_{x}\gamma(x)-\frac{1}{2}}n^{1-\zeta_{x}\gamma(x)}\right),$
which vanishes as $n\to\infty$. ∎
@@
As above, it holds that $\overline{F}^{x}(tu_{n}^{x})/\overline{F}^{x}(u_{n}^{x})=\overline{F}^{x}(tu_{n}^{x})n/k_{n}$. Consequently
$\left|\frac{\overline{F}^{x}(tu_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}-t^{-1/\gamma(x)}\right|$
$=\frac{n}{k_{n}}\left|\overline{F}^{x}(tu_{n}^{x})-c_{x}\log(u_{n}^{x}t)(u_{n}^{x}t)^{-1/\gamma(x)}+c_{x}\log(u_{n}^{x}t)(u_{n}^{x}t)^{-1/\gamma(x)}-\overline{F}^{x}(u_{n}^{x})t^{-1/\gamma(x)}\right|$
$\leq\frac{n}{k_{n}}\Big(\left|\overline{F}^{x}(tu_{n}^{x})-c_{x}\log(u_{n}^{x}t)(u_{n}^{x}t)^{-1/\gamma(x)}\right|$
$\qquad+t^{-1/\gamma(x)}\left|c_{x}\log(u_{n}^{x})(u_{n}^{x})^{-1/\gamma(x)}-\overline{F}^{x}(u_{n}^{x})\right|+c_{x}\log(t)(u_{n}^{x}t)^{-1/\gamma(x)}\Big)$
$\leq\text{cst}\frac{n}{k_{n}}\left((tu_{n}^{x})^{-\zeta_{x}}+t^{-1/\gamma(x)}(u_{n}^{x})^{-\zeta_{x}}+c_{x}\log(t)(u_{n}^{x}t)^{-1/\gamma(x)}\right)$
$\leq\text{cst}\frac{n}{k_{n}}t^{-1/\gamma(x)}(u_{n}^{x})^{-1/\gamma(x)}\left(1+\log(t)\right).$
@@
$(u_{n}^{x})^{-1/\gamma(x)}\log(u_{n}^{x})=\frac{k_{n}}{n}+\mathcal{O}\left((u_{n}^{x})^{-\zeta_{x}}\right),$